if a fish is moving in the upward directon with speed 2m/s and a bird is moving in downward direction with 3m/s , then find the velocity of bird with respect to the fish .. (refractive index of water wrt air is 4/3)
 
 the ans is 6m/s but plx explain me how to solve these type of questions ...

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

In this question, observe that we have to find speed of ‘bird with respect to fish’. The fish always sees the bird at a height greater than actual height and so it will also observe speed to be greater than actual speed.

Apparant heighth'=\mu h
\impliesApparent speed=\dfrac{dh'}{dt}=\mu \dfrac{dh}{dt}
Hence apparent speed of bird with respect to fish would be\dfrac{4}{3}\times 3m/sdownwards=4m/sdownwards.

Hence velocity of bird with respect to fish=\vec{v_{bird}}-\vec{v_{fish}}=4m/s+2m/s=6m/sdownwards.